OpenStudy (anonymous):

If each side s of a square is increased by 1%, then the area is increased approximately: 0.02s 0.02s^2 0.01s^2 1% 0.01s I dont get this one either

OpenStudy (anonymous):

sorry but can you help?

OpenStudy (the_fizicx99):

0.01s^2 because its a square ?

OpenStudy (anonymous):

its an AP calc question i dont think thats it

OpenStudy (anonymous):

area = s^2 new area = [s(1.01])^2 = (s^2)(1.01)^2 difference is ... I am not supposed to give answers.

OpenStudy (loser66):

@douglaswinslowcooper don't worry, he is an AP cal student

OpenStudy (anonymous):

oh i see kinda what you did

OpenStudy (anonymous):

so its 0.02s^2

OpenStudy (mathmale):

I think you're dealing with "differentials" here. If the area of a square of side s is A=s^2, then the derivative w. r. t. s is dA/ds = 2s, and the resulting differential is dA=2sds. Supposing that the initial side of the square was s and the increase in s is 1%, then the increase in s can be expressed as 0.01s. Then dA = 2 s ds = 2 s (0.01s). Simplify this expression. That's the approx. increase in the area of the square.

OpenStudy (anonymous):

Yes, 2%. In general, for small percentage changes if f(x) = x^n, a p percentage change in x will produce an np percentage change in f. 1% change in length produces 3% change in volume etc.

OpenStudy (kagıtucak):

take length as 100 then %1 of it is 1 then new lentgh become 101 at the beginnnig area is 10.000 then it becomes 10.201 so there is a 201 m2 increasing so 201/10.000*100 =2.01

OpenStudy (anonymous):

yes this is a differentials question thank you @douglaswinslowcooper